OpenStudy (anonymous):

Part 1: Create your own factorable trinomial. Part 2: Explain, in complete sentences, how the trinomial is factored. Part 3: Explain, in complete sentences, the process used to check the factors for accuracy.

OpenStudy (saifoo.khan):

Part 1: x^2 - 5x + 6

OpenStudy (saifoo.khan):

(x - 3)(x - 2) But you can always apply the formula for the roots of a quadratic: ax² + bx + c [-b +/- √(b² - 4ac)]/2a In your case its [5 +/- √(25 -4 * 6 * 1] / 2 = (5 +/- √1)/2 = 6/2 and 4/2 = 3 and 2 So your roots are (x - 3) and (x - 2)